OpenStudy (anonymous):

How do I solve a question like : 1.1k=1.2K-5.4=-10 Show your work using inverse operations P.S will give a medal for best answer :)

OpenStudy (anonymous):

is it small K and k?

OpenStudy (anonymous):

sorry just k

OpenStudy (anonymous):

isnt it just 1.1k=1.2k-5.4 ? why need -10? 0.1k=5.4 k= 54

OpenStudy (anonymous):

was wondering if u typed the question correctly.. mind checking again?

OpenStudy (anonymous):

no the equation is the one i posted you have to solve for k :)

OpenStudy (anonymous):

1.1k=1.2K-5.4 so k = (1.2K=5.4)/1.1 then ur second equation 1.1k=-10 so k = -10/1.1

OpenStudy (anonymous):

lol the 1.1k is useless?

OpenStudy (anonymous):

uhhh i am really confused...

OpenStudy (theviper):

These question is wrong :P

OpenStudy (anonymous):

It is in my textbook so I dont know...

OpenStudy (theviper):

But sometimes books are also wrong ;)

OpenStudy (anonymous):

ok FAANNKSS :)

OpenStudy (theviper):

:)

OpenStudy (theviper):

U deserve a medal in making a book wrong :D

OpenStudy (anonymous):

:)

OpenStudy (theviper):

:)